# Generating Chain-of-Thoughts with a Direct Pairwise-Comparison Approach to Searching for the Most Promising Intermediate Thought

### v1

@@ C Cost and Efficiency
In Table 4, we report the token cost and average accuracy comparison of our proposed approaches with other contenders. We can observe that the token costs of our proposed C-ToT approaches and the contenders are task-specific and generally incomparable. Since we preserve all previous intermediate thoughts in the QA task, the token cost of the C-ToT approaches is higher than that of the S-ToT algorithm, but the performance is simultaneously improved. In the QA and Game of 24, the token costs of the Comp-SToT and S-ToT approaches are comparable, but the Comp-SToT approaches achieve better average accuracy. In the Sudoku puzzle, the token cost of Comp-SToT is lower than that of S-ToT. These results indicate the effectiveness of using the direct pairwise comparison approach to find the most promising intermediate thoughts. In practice, we can further reduce the token cost by using a counter for each intermediate thought to track its comparison frequency. If the comparison count of an intermediate thought exceeds a threshold, we can remove it from the tree.
@@
Table 4: Average Accuracy of different methods against their token costs.
Method Question Answering Game of 24 Sudoku Puzzle
Generate/Prompt tokens Cost per case Accuracy Generate/Prompt tokens Cost per case Accuracy Generate/Prompt tokens Cost per case Accuracy
CoT 106/136 0.0003 42.3% 99/437 0.0006 4.3% 431/178 0.001 44.4%
SC-CoT 1647/2023 0.0054 58.4% 1717/6555 0.010 8.0% 6292/2666 0.015 47.8%
SToT 1551/5415 0.0085 57.1% 1368/12205 0.015 34.3% 6309/23933 0.037 60.0%
SC-SToT 2081/13459 0.018 57.6% 2284/40825 0.045 40.0% 6568/70129 0.083 66.7%
Comp-SToT 1515/6299 0.0093 59.0% 1309/11963 0.015 36.3% 2666/13164 0.019 65.6%
Back-SToT 1551/8135 0.011 58.0% 1679/23178 0.027 39.0% 4536/21383 0.030 72.2%
C-ToT (Stand.) 1498/14627 0.017 61.4% 2452/21003 0.026 40.0% 5340/29565 0.040 74.4%
C-ToT (Duel.) 1649/52044 0.055 63.0% 2174/60578 0.065 41.0% 7148/86425 0.101 75.5%

### v2

@@ C.2 Cost and Efficiency
We can observe that the token costs of our proposed C-ToT approaches and the contenders are task-specific and generally incomparable. Since we preserve all previous intermediate thoughts in the QA task, the token cost of the C-ToT approaches is higher than that of the S-ToT algorithm, but the performance is simultaneously improved. In the QA and Game of 24, the token costs of the Comp-SToT and S-ToT approaches are comparable, but the Comp-SToT approaches achieve better average accuracy. In the Sudoku puzzle, the token cost of Comp-SToT is lower than that of S-ToT. These results indicate the effectiveness of using the direct pairwise comparison approach to find the most promising intermediate thoughts. In practice, we can further reduce the token cost by using a counter for each intermediate thought to track its comparison frequency. If the comparison count of an intermediate thought exceeds a threshold, we can remove it from the tree.
Method Generate/Prompt tokens Cost per case Accuracy
CoT 106/136 0.0003 42.3%
SC-CoT 1647/2023 0.0054 58.4%
SToT 1551/5415 0.0085 57.1%
SC-SToT 2081/13459 0.018 57.6%
Comp-SToT 1515/6299 0.0093 59.0%
Back-SToT 1551/8135 0.011 58.0%
C-ToT (Stand.) 1498/14627 0.017 61.4%
C-ToT (Duel.) 1649/52044 0.055 63.0%
@@
Table 6: Average accuracy of different methods with token costs on QA.
Method Generate/Prompt tokens Cost per case Accuracy
CoT 99/437 0.0006 4.3%
SC-CoT 1717/6555 0.010 8.0%
SToT 1368/12205 0.015 34.3%
SC-SToT 2284/40825 0.045 40.0%
Comp-SToT 1309/11963 0.015 36.3%
Back-SToT 1679/23178 0.027 39.0%
C-ToT (Stand.) 2452/21003 0.026 40.0%
C-ToT (Duel.) 2174/60578 0.065 41.0%
Table 7: Average accuracy of different methods with token costs on Game of 24.
Method Generate/Prompt tokens Cost per case Accuracy
CoT 431/178 0.001 44.4%
SC-CoT 6292/2666 0.015 47.8%
SToT 6309/23933 0.037 60.0%
SC-SToT 6568/70129 0.083 66.7%
Comp-SToT 2666/13164 0.019 65.6%
Back-SToT 4536/21383 0.030 72.2%
C-ToT (Stand.) 5340/29565 0.040 74.4%
C-ToT (Duel.) 7148/86425 0.101 75.5%
Table 8: Average accuracy of different methods with token costs on Sudoku.
C.3 Ablation Studies
We report the ablation studies on Aqua dataset, while we also observe the same trend in other datasets. 5 test runs were conducted and the average accuracy and the number of tokens (completion tokens / prompt tokens) are reported.
We study the number of intermediate thoughts generated and selected in each round and report the results in Table 9. With a fixed number of thoughts generated each round, selecting more thoughts leads to higher costs, while accuracy may not benefit much. With a fixed number of thoughts selected each round, generating more thoughts leads to a significant increase in accuracy because we can explore more thoughts. These results could benefit the further use of CoT methods.
Generate $m$ / Select $K$ $K=1$ $K=2$ $K=3$ $K=5$ $K=6$
$m=1$ 41.6% ; 50/515
$m=3$ 46.7% ; 139/1525 50.2% ; 244/2565 49.8% ; 375/3658
$m=5$ 50.1% ; 288/2653 53.5% ; 495/4452 55.0% ; 549/5945 54.8% ; 882/9379
$m=10$ 53.0% ; 518/5189 56.6% ; 896/8716 57.3% ; 1091/11875 57.5% ; 2120/19473 57.6% ; 2155/20982
$m=12$ 53.8% ; 674/6334 57.4% ; 1212/10732 61.4% ; 1498/14627 61.4% ; 1531/16002 61.5% ; 2125/26329
Table 9: Ablation studies on parameters.
We study the threshold for removing intermediate thoughts vs. accuracy and cost and report the results in Table 10. Under the same tree depth, different thresholds show relatively stable performance.
Depth $d$ / Threshold Th Th$=1$ Th$=2$ Th$=3$ Th$=4$ Th$=5$
$d=3$ 58.0% ; 1062/11337 60.7% ; 1315/13611 61.0% ; 1478/15111 61.4% ; 1513/15713 61.4% ; 1558/16107
$d=4$ 58.8% ; 1426/13792 60.9% ; 1732/18822 61.4% ; 1771/19940 61.6% ; 1804/20003 61.3% ; 2001/20312
$d=5$ 60.0% ; 1760/17556 61.0% ; 2143/21006 61.6% ; 2271/23412 61.7% ; 2265/24031 61.3% ; 2425/24755
Table 10: Ablation studies on the number of thresholds for removing intermediate thoughts.
We study the number of comparison and report the results in Table 11. In general, increasing the number of comparison results in better accuracy and higher cost.
Dataset / Comparison $n$ $n=1$ $n=3$ $n=5$ $n=8$
Aqua 61.4% ; 1498/14627 63.0% ; 1649/52044 64.3% ; 1824/78194 64.7% ; 1997/128802
Table 11: Ablation studies on the number of comparisons.